### v12

@@ 8 Optimum investment policy
Data and arguments suggest that the optimum amount of saving, at the global scale, is replacement saving to offset depreciation, and nothing more. That would not mean book depreciation, as this study has stressed differences between book and market values. Up to a point, it should be possible to analyze the composition of market capital, and to model depreciation of the whole. A better plan, as Solow (1956) wrote in response to Harrod’s knife edge argument (Harrod, 1939), is to trust the market to maximize rate of return, and to sense the point where glut begins and returns fall. [^9] Harrod had argued that saving must hit the warranted rate exactly or risk positive feedback through the operation of the output/capital ratio (accelerator). Markets do so imperfectly when tax and other public policy reward saving over distributions and consumption. Findings in this paper suggest review of such policies. These include the double tax on dividends, and the greater tax rate on ordinary income than on capital gains. Effects of removing the double tax, and removing the difference between tax rates on ordinary income and on capital gains, could be revenue-neutral and non-partisan if the corporate tax were raised to match, if the tax rates on ordinary income and on capital gains met somewhere between, and if thoughtful grandfathering eased the transition.
@@ 9 Optimum government policy
Households dissave in consumption, while firms dissave through dividends and other distributions. Public policy, in many nations, has inhibited dissaving in the belief that such policies favor capital growth. Rather they inhibit consumption, add nothing to capital growth, and thus inhibit net output as a function of both together (see Appendix B).
The effect of repealing such policies would be to address capital glut not by investing less, but by disinvesting more. [^10] Higher dividends and distributions buy firms, that is, would enable higher consumption by households. This strategy spares government the burden of guiding investment, at least until the effects of the repeal are known.
10 Inferences for national accounts and macroeconomics
We have attempted to show that net saving does not tend to equate to capital growth, and indeed contributes no part of capital growth. If our argument proves correct, then national accounts, and the teachings of macroeconomics (macro) which they express, are constructed on a fundamental misperception. We would therefore welcome any critiques of our findings as a matter of academic and public interest.
@@ 11 Inferences for microeconomics
Microeconomics ("micro") studies the mutual effects of supply, demand and price. It teaches that producer cost, including interest as the cost of time, tends to equilibrate to market value realized. We have proposed an exception in the innovator’s reserve; the innovator pays the going rates for skills and materials needed, but applies them to make products of greater market value.
12 Inferences for inflation
If free growth theory is correct, then net investment adds a superfluous cost to capital growth, while decreasing market value per unit through capital glut. Both effects would tend to contribute directly to capital inflation, and indirectly to inflation of consumer goods by adding depreciation costs. This consideration adds to the case for urgent attention to our argument, and for more study to confirm or refute it.
13 Net output
In Keynesian tradition, net output equals consumption plus net investment. We have disputed this reasoning in our argument that net investment adds nothing to capital growth, inhibits consumption, and thus tends to reduce the sum of both. We also critique it on another ground. Assume, for the sake of argument, that we are wrong, and that net investment realizes equal capital growth. Net output means value added. Appendix B will show an argument that to interpret value added as the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative (human depreciation) and one positive (self-invested work), in value added to human capital. Both of these flows are unmeasurable by known means. Thus we have not sought to measure net output or its growth, and have treated capital growth alone.
@@ 14 Data sources
All our data are drawn from the free online website World Inequality Database (WID). This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations. We show results for the 92 of those countries which report both factors, net saving and market-value capital, needed for deriving the thrift indexes. WID’s source for these data is national accounts.
@@ 15 Accessing our results and methods
16 Displays
The web appendix includes displays of $\theta$ for each year in each country over the report period. These tend to show upward and downward spikes in values of $\theta$ in some years. Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years. Small denominators magnify errors in measurements of numerators. Worse, when denominators are small, small mismeasurements in them might reverse the term in sign.
To maximize reliability of test results, we apply a range of screens to omit years where absolute denominators fall below a given threshold. Some displays in the web appendix show data for all years, regardless of denominator size. Others screen out all years where absolute denominators are less then 0.01, then 0.025, then continuing upward in increments of 0.025 to a maximum screen of 0.15. $\theta$ is plotted for each country unscreened and at each of the seven successive levels of screening. All displays in the main body of this paper apply a screen of 0.01. The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
@@ 17 Disclaimers
Evidence supports the proposal of Modigliani (1954) that the young save to support retirement, and that retirees dissave to support consumption. [^11] Friedman (1957) follows a different argument to a similar conclusion. Such age-based differences are expected to offset in the stationary state, where tradition and free growth theory agree that overall net saving is zero. We accept that individual net saving can realize equal individual capital growth, so long as collective saving does not exceed collective dissaving.
@@ 18 Discussion and conclusions
Capital glut is the condition warned against by West, Ricardo, Malthus and Harrod. It is loosely defined as oversupply of capital at the current state of technology. We take it to mean or include a state where more investment does not realize more capital value. Findings shown in our displays suggest that net saving raises the physical quantity of capital, say in number of shops, manufacturing plants or finished goods of similar design, without raising aggregate value of capital, and so contributes to capital glut by that understanding.
These findings challenge the teachings that capital growth is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the opportunity cost of capital, converges to market realization. Evidence suggests that all capital growth is free, and consequently that market realization, in the presence of innovation, exceeds producer cost by the entirety of capital growth. Meanwhile the same evidence, which indicates that net saving adds no capital value, suggests a review of the teachings that consumption plus net saving gives net income, and that consumption plus net investment gives net output. Appendix B will also question the latter teachings.
Embodied growth is disembodied growth on a finer scale. It redeploys or repurposes existing labor skills, raw materials, and plant capacity, as well as existing finished goods, to achieve higher returns than available from the customary uses which determine their prices. The present value of yields from this advantage in return, or equivalently the innovator’s reserve, defines the non-random component in free growth.
Solow’s puzzle runs deep. How can thoughts be things? How can assessments of present value pay, at least in part, for their own realization? Science and philosophy push our minds into areas where they are not at home. Neither relativity nor quantum mechanics, which explore the scale of great speeds and minute distances, sink in easily. Economics, too, invites surprises as it looks farther from the scale of the household and community and firm to which we are accustomed. Free growth theory predicts only at the universal scale, and accepts that thrift explains a greater part of growth at progressively smaller scales. Our findings, with that allowance, are awkward for much of economics as generally taught. We encourage others to check our data and methods and inferences, and improve on them wherever they can.
@@ Appendix A.  Equivalence of saving and investment
Keynes (1936) argued correctly that saving, if not invested in currently produced producer goods and services, adds nothing to capital, and thus nothing to output. He inferred that uninvested saving does not achieve actual saving, as it accomplishes no growth, and described it as "intended saving." True saving, by that reasoning, equalled investment.
We suggest that investors seek to maximize return, within risk tolerance, and will sometimes hold saving in cash or equivalents at zero return in recessions and depressions when positive returns cannot be found, and so when investing would tend to reduce capital and output rather than increase them. Thus we do not share the Keynesian view that saving need be invested in order to maximize its value. Saving will tend to equal investment in normal times, when prospective returns bring animal spirits, and could include saving in cash when not. It is only for these profitable times that we accept the equality of saving and investment, and only as a simplifying assumption. [^12] The weakness in the Keynes argument, if we are right, was his inference that saving does not accomplish its purpose unless invested to realize capital growth. Its purpose, we argue, is to maximize capital growth according to investment opportunities, and to maximize it by withholding investment when all investments promise loss.
@@ Appendix B. Net output with human capital
It is possible in principle to model a value-added chain which includes human capital, and to compare findings with those shown in Eq. (B.1). Let human capital $H$, in that new model, stand as the last link in the value-added chain. Adapting the classic illustration of the value added principle, say that farms produce wheat, mills convert the wheat to flour, bakeries convert the flour into bread, and humans convert some of the bread, called invested consumption, into human capital. The net outflow from this extended value-added chain is not all of consumption, but only pure consumption, meaning the part remaining after the part invested in human capital (invested consumption) is subtracted. [^13] The concepts of invested consumption, pure consumption, self-invested work and human depreciation were introduced in Schultz (1961). By this reasoning, the principle that net output is expressed in capital growth plus net outflow gives
@@
Yoram Ben-Porath (1967) reasoned that growth in human capital equals invested consumption plus self-invested work less human depreciation [^14] Equation 4 in Ben-Porath’s paper, summarizing his first three equations. His terms and notation differ from ours.. Let $C_{s},\ W_{s}$ and $D(H)$ show these flows respectively. Thus the combined arguments of Schultz and Ben-Porath arrive at
@@ Appendix C. Mill’s statement of the free growth idea
This passage may be the first clear statement of what we call free growth theory. Mill’s use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours, and that he meant what we call free growth rather than the alternating phases of higher and lower saving rates pictured in thrift theory.

### v13

@@ 9 Optimum investment policy
Data and arguments suggest that the optimum amount of saving, at the global scale, is replacement saving to offset depreciation, and nothing more. That would not mean depreciation found by the depreciated cost method, which equates net investment to capital growth at all scales, and which cannot measure the contributions of the innovator’s reserve or tech externalities. Up to a point, it should be possible to analyze the composition of market-value capital, and to model depreciation of the whole, just as national balance sheet accounting does with replacement value capital today. A better plan, as Solow (1956) wrote in response to Harrod’s knife edge argument (Harrod, 1939), is to expect the market to maximize rate of return, and to sense the point where glut begins and returns fall. [^15] Harrod had argued that saving must hit the warranted rate exactly or risk positive feedback through the operation of the output/capital ratio (accelerator). Markets do so imperfectly when tax and other public policy reward saving over distributions and consumption. Findings in this paper suggest reconsideration of such policies. These include the double tax on dividends, and the greater tax rate on ordinary income than on capital gains. Effects of removing the double tax, and removing the difference between tax rates on ordinary income and on capital gains, could be revenue-neutral and non-partisan if the corporate tax were raised to match, if the tax rates on ordinary income and on capital gains met somewhere between, and if thoughtful grandfathering eased the transition.
@@ 10 Optimum government policy
Households dissave in consumption, while firms dissave through dividends and other distributions. Public policy, in many nations, has inhibited dissaving in the belief that policies to promote net saving favor capital growth. Rather they inhibit consumption, add less than 8% to capital growth, and thus inhibit net output as a function of both together (see Appendix B).
The effect of repealing such policies would be to address capital glut not by investing less, but by disinvesting more. [^16] Higher dividends and distributions by firms, that is, would enable higher consumption by households. Demand for dividends should rise in response as shareholders are freed from the burden of the double tax, and firms would be incentivized to compete to meet that demand. As firms did so, and raised dividends, firms would find a lower share of income left to invest, and would tend to invest more selectively. Lower-return investment opportunities would tend to find fewer bidders, and average returns, in consequence, should tend to rise. Government might watch and wait, in such a scenario, and weigh other policy as outcomes appear.
@@ 11 Inferences for microeconomics
Microeconomics ("micro") studies the mutual effects of supply, demand and price. It teaches that producer cost, including interest as the cost of time, tends to equilibrate to market value realized. We have proposed an exception in the innovator’s reserve; the innovator pays the going rates for skills and materials needed, and the going interest rate as the price of time, but applies them to make products of greater market value (See Appendix D).
12 Inferences for interpretation of national accounts
National accounts define the sum of consumption and investment as product. Common teaching identifies this sum, in its version net of depreciation, as net output in the sense of value added. We have disputed this reasoning in our argument that net investment realizes less than 8% of its cost in capital growth, inhibits consumption by its full cost, and thus tends to reduce the sum of both. We also critique it on another ground. Assume, for the sake of argument, that we are wrong, and that net investment realizes equal capital growth. Appendix B will show an argument that to interpret value added as the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative (human depreciation) and one positive (self-invested work), in value added to human capital. Both of these flows are unmeasurable by known means. Thus we have not sought to measure net output or its growth, and have treated capital growth alone.
Common teaching also equates the sum of investment and consumption to expenditure. We agree that spending on new goods and services can find no other dispositions, and that investment and consumption can find no other sources, if we recognize imputed payment for voluntary consumption services as in parenting.
By the argument in Appendix B noted above, consumption and market-value capital growth give the measurable components of true net output in the sense of value added. Then their sum, which might be thought of as market-value net product, would give a measurable proxy for net output, although not net output per se.
It appears from the above that gross product and what we call market-value net product are both informative. The first gives expenditure, although not value added, and the second gives at least a first approximation of value added.
@@ 13 Data sources
All our data are drawn from the free website World Inequality Database wid.world.com (WID). This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations. We show results for the 92 of those countries which report both factors, net saving and market-value capital, needed for deriving the thrift indexes. WID’s source for these data is national accounts.
@@ 14 Accessing our results and methods
15 Screening out small denominators
Our displays include findings for $\theta$ for each year in each country over the report period. These tend to show upward and downward spikes in values of $\theta$ in some years. Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years. Small denominators, even if measured perfectly, magnify errors in measurements of numerators. Worse, when denominators are small, small mismeasurements in them might reverse the term in sign.
To maximize reliability of test results, we apply a range of screens to omit years where absolute denominators fall below a given threshold. Some displays in the web appendix show data for all years, regardless of denominator size. Others screen out all years where absolute denominators are less then 0.01, then 0.025, then continuing upward in increments of 0.025 to a maximum screen of 0.15. $\theta$ is plotted for each country unscreened and at each of the seven successive levels of screening. The screen chosen in displays in this paper is 0.01. The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
@@ 16 Disclaimers
Evidence supports the proposal of Modigliani (1954) that the young save to afford families and retirement, and that retirees dissave to pay for both as they arrive. [^17] Friedman (1957) follows a different argument to a similar conclusion. Such age-based differences are expected to offset in the stationary state, where tradition and free growth theory agree that overall net saving is zero. We accept that individual net saving can realize equal individual capital growth, so long as collective saving does not exceed collective dissaving.
@@ 17 Discussion and conclusions
With disclaimers noted, and with allowance for uncertainties of the thrift residue, we have argued that the innovator’s reserve and tech externalities pay for capital growth as a whole, and that net saving by some can realize equal capital growth only insofar as others dissave as much. Market appraisals raise present value at the warranted rate with no need for net saving at the overall scale. We adopted the Modigliani model, where young households save to afford families and retirement while older households dissave to pay for both as they arrive, as an example of this saving/dissaving balance.
These findings challenge the teachings that capital growth at all scales is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the cost of time, converges to market realization. Evidence suggests that much or all capital growth at the collective scale is free, and consequently that market value, in the presence of innovation, exceeds producer cost by at least the majority of capital growth. Meanwhile the same evidence suggests a review of the teachings that consumption plus net saving gives net income, and that consumption plus net investment gives net output. Appendix B will also question the latter teachings.
Embodied growth is disembodied growth on a finer scale. It redeploys or repurposes existing labor skills, raw materials, and plant capacity, as well as existing finished goods, to achieve higher returns than available from the customary uses which determine their prices. The present value of yields from this advantage in return, or equivalently the innovator’s reserve, plus tech externalities, defines the non-random component in free growth.
Solow’s puzzle runs deep. How can thoughts be things? How can assessments of cash flows over the future pay for their realization in the present, even when plant of new design must come first? Science and philosophy push our minds into areas where they are not at home. Neither relativity nor quantum mechanics, which explore the scale of great speeds and minute distances, sink in easily. Economics, too, invites surprises as it looks farther from the scale of the household and community and firm to which we are accustomed. Free growth theory predicts only at the universal scale, and accepts that thrift explains a greater part of growth at progressively smaller scales. Our findings, with that allowance, are awkward for much of economics as generally taught.
The data are compelling, even so, and our explanatory hypothesis in Section 1 fits common observation. We can see that producing more from the same inputs can spare a need for cutting back elsewhere. And we know that markets invest according to forecast future earnings, even when present earnings are nil. Free growth theory adds only that these well-understood effects, described in Piketty and Zucman (2014) and Kurz (2023), explain all or substantially all of capital growth.
@@ Appendix A.  Equivalence of saving and investment
Keynes (1936) argued correctly that saving, if not invested in currently produced producer goods and services, adds nothing to capital, and thus nothing to output. He inferred that uninvested saving has no value, as it accomplishes no growth, and described it as "intended saving." True saving, by that reasoning, equalled saving actually invested.
We suggest that investors seek to maximize return, within risk tolerance, and will sometimes hold saving in cash or equivalents at zero return in recessions and depressions when positive returns cannot be found, and so when investing would tend to reduce capital and output rather than increase them. Thus we do not share the Keynesian view that saving need be invested in order to maximize its value. Saving will tend to equal investment in normal times, when prospective returns bring animal spirits, and could include saving in cash when not. It is only for these profitable times that we accept the equality of saving and investment, and only as a simplifying assumption. [^18] The weakness in the Keynes argument, if we are right, was his inference that net saving finds no utility, and hence no economic value, except as invested to realize capital growth. Its utility in a more general sense, we suggest, is to maximize capital growth according to investment opportunities, and to maximize it by withholding investment when all investments promise loss.
@@ Appendix B. Net output with human capital
It is possible in principle to model a value-added chain which includes human capital, and to compare findings with those shown in Eq. (B.1). Let human capital $H$, in that new model, stand as the last link in the value-added chain. Adapting the classic illustration of the value added principle, that is, where farms produce wheat, mills convert the wheat to flour, and bakeries convert the flour into bread, say that humans convert some of the bread, called invested consumption, into human capital. The net outflow from this extended value-added chain is not all of consumption, but only pure consumption, meaning the part remaining after the part invested in human capital (invested consumption) is subtracted. [^19] The concepts of invested consumption, pure consumption, self-invested work and human depreciation were introduced in Schultz (1961). By this reasoning, the principle that net output is expressed in capital growth plus net outflow gives
@@
Yoram Ben-Porath (1967) reasoned that growth in human capital equals invested consumption plus self-invested work less human depreciation [^20] Equation 4 in Ben-Porath’s paper, summarizing his first three equations. His terms and notation differ from ours. He expressed his idea in Cobb-Douglas functions, whose values vary with scalars and exponents chosen, as opposed to our simpler algebra. We believe, however, that meanings agree.. Let $C_{s},\ W_{s}$ and $D(H)$ show these flows respectively. Thus the combined arguments of Schultz and Ben-Porath arrive at
@@ Appendix C. Mill’s statement of the free growth idea
This passage may be the first clear statement of what we call free growth theory. Mill’s use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours.
Appendix D. Present value and cost
We use the term cash flow to mean value transferred from an asset to its owner, in cash or in kind, less value transferred from the owner to the asset. Thus cash flow will generally be negative over any period of investment by owners in buying or building the asset, and then positive when investment ends and net yields begin. Present value of the asset, at any moment, is the sum of future cash flows expected from that moment on, each discounted at the market time discount rate (see Appendix F). Thus the present value of a differential amount of future cash flow expected at time $t$, shown as $dF(t)$, if discounted at the market rate $r_{m}$, finds a present value $dPV$ given by
$dF(t)e^{-r_{m}t}=dPV\ .$ (D.1)
Present value of all cash flows together, then, is
$PV=\int^{\infty}F(t)e^{-r_{m}t}dt\ ,$ (D.2)
where the implicit lower limit of integration is the present time zero.
The actual investment needed to realize this present value will vary inversely with the productivity of the investment. Productivity of a given investment is measured in the rate of total return $r_{I}$ it generates, composed of growth rate $g(K)$ and cash flow rate. Consider a differential amount of investment $dI$ invested now, and yielded out wholly, including both principal and accumulated growth, in a differential amount of cash flow $dF$ at a single later time tt.212121dIdI and $dF$ need not represent the whole of investment or cash flow at times $t_{1}$ and $t_{2}$. It is always possible, in principle, to divide a given investment, including appreciation, into infinitesimal (differential) parts each yielded out at a single later moment. As no part of principal or accumulated growth from this particular investment was yielded during the interim $t$, by assumption, growth rate $g(K)$, over that interim, equalled the whole of rate of return $r_{I}$. Hence
$dF(t)=dIe^{r_{I}t}\ ,\quad\text{or equivalently}\quad dI=dF(t)e^{-r_{I}t}\ .$ (D.3)
Investment $I$ is the sum of these differential components $dI$. That is,
$I=\int^{\infty}dI=\int^{\infty}dF(t)e^{-r_{I}t}dt\ .$ (D.4)
Consider the case where $r_{I}$ and $r_{m}$ are equal. By Eqs. (D.1) through (D.4),
$dPV=dFe^{-r_{m}t}=dFe^{-r_{I}t}=dI\quad\text{and}\quad PV=I\ ,\quad\text{if}\quad r_{I}=r_{m}\ .$ (D.5)
Now relax that assumption, and allow $r_{m}$ and $r_{I}$ to differ. The innovator’s reserve is defined as present value less cost invested. Thus the component in each differential cash flow due to the innovator’s reserve is
$dPV-dI=dF(t)e^{-r_{m}t}-dF(t)e^{-r_{I}t}=dF(t)(e^{-r_{m}t}-e^{-r_{I}t})\ .$
The innovator’s reserve as a stock, which we show as $RSV$, is accordingly found at
$RSV=PV-I=\int^{\infty}F(t)e^{-r_{m}t}dt-\int^{\infty}F(t)e^{-r_{I}t}dt=\int^{\infty}F(t)(e^{-r_{m}t}-e^{-r_{I}t})dt\ .$ (D.6)
These equations confirm that present value equals investment needed when $r_{I}$ equals $r_{m}$, and exceeds investment needed when $r_{I}$ exceeds $r_{m}$.
Appendix E. Analysis of the thrift residue
The data shown in our displays find the thrift residue by three measures ranging from 7.71% to 5.59%. This residue gives the extent to which $s^{*}$ and $g(K)$ are causally related, but does not give the direction of causality. It shows the extent to which change in $s^{*}$ brought about a change in $g(K)$, or conversely, or both were caused primarily by a common third factor. In the last case, the third factor might bring about a direct effect on $s^{*}$ or $g(K)$ alone, at first, which might then cause a reaction by the other. Thrift theory predicts that a change in $s^{*}$ will bring an equal change in $g(K)$ before allowance for market noise, whether the change in $s^{*}$ was itself the primary cause, or whether it was a proximate cause driven by an exogenous shock, say war or epidemic or other disruption in supply or demand, as the primary one.
Internal misjudgement can contribute to these setbacks by weakening defences, or can bring them about without need for significant outside causes. Some economists attribute at least some recessions to underspending on currently produced consumer and producer goods, made worse by stock market skittishness. This interpretation would fit thrift theory in picturing a shortfall in net investment as causing the drop in market value of capital.
Another view pictures recessions as caused or facilitated by accumulated overinvestment through overestimation of growth prospects, and pictures recession at the end as a reality check. That scenario would reverse the direction of causality expected in thrift theory by putting the drop in $g(K)$ and income first, and by seeing the drop in $s^{*}$, the ratio of net saving to capital, as driven by a need to maintain consumption as income falls. The same chain of causality would repeat in the opposite direction as recovery from these crises brings capital value and income up as a direct effect, and so enables households to save against the next crisis.
This issue was debated famously by Keynes and the Austrian school during the great depression, by Malthus, Say and Ricardo two centuries ago, and Smith and Mandeville before. They did not have the data reported in our displays. We suggest that the thrift theory explanation might have been urged less confidently in the light of information that net investment has accounted for less than 8% of capital growth.
Other points can be added. Why should investors trade less efficiently in market equities than in the goods bought and sold by the underlying firms? This is the question we posed in our discussion of Tobin’s $q$ in Section 8.
Appendix F. Clarifying cash flow and $r_{m}$
The term "cash flow" is not much used in economics outside the specialty of finance, where it is taken as the flow discounted to find present value, or added to capital appreciation to find total return. We extend this concept to non-financial as well as financial assets by interpreting the services or utilities for which we buy a product as cash flows, or net withdrawals of value from assets to owners. [^22] We buy a non-financial asset, say a book or chair or car or house, for the services (taste satisfactions) we expect from it, each discounted at the risk-adjusted time discount rate we each apply. As each expected satisfaction eventuates, we deduct its value from the pool of remaining expected satisfactions, and so from remaining value to us of the asset. Although we do these calculations intuitively rather than formally, they fit the essential criteria for of finding present value as the sum of time-discounted projected cash flows. These cash flows are discounted to present value at the risk-adjusted time preference rate which the market applies to each product. Free growth theory predicts only at the universal scale, and is interested primarily in cash flows and time-discount rates at the scale of all products and all capital together. Thus we define $r_{m}$, except where stated otherwise, as the market time-discount rate, or equivalently the cap-weighted average of risk-adjusted rates, for the economy overall.
Appendix D argued that $r_{m}$ also gives the market-wide rate of return in the absence of innovation, but that productivity gain through innovation can bring rate of return above the time discount rate. Thus $r_{m}$ can also be understood as the current standard rate of return before allowance for innovation. Note also that diminishing returns brought by capital glut should tend to drive $r_{m}$ down, as the measure of overall return before effects of current innovation, and so make further investment easier to justify by enhancing present value under Eq. (D.2). In this sense, capital glut may tend to be self-perpetuating.